OpenStudy (anonymous):

Factor the following p^3-125q^3

OpenStudy (anonymous):

p^3-(5q)^3

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

Oh is this a difference of cubes?

OpenStudy (anonymous):

HINT: a^3 - b^3 =(a-b)(a^2+2ab+b^2)

OpenStudy (anonymous):

Ah ok got it. I will tell you in a second.

OpenStudy (anonymous):

I got (p-5)(p^2+10p+25)

OpenStudy (anonymous):

u forgot q

OpenStudy (anonymous):

Something is missing

OpenStudy (anonymous):

(p-5q)(P^2-10pq+25q^2

OpenStudy (anonymous):

)*

OpenStudy (anonymous):

Is this right?

OpenStudy (anonymous):

(p-5q)(p^2+5pq+25q^2)

OpenStudy (anonymous):

Sorry gave the wrong formula: Actually, a^3-b^3=(a-b) (a^2+ab+b^2)